OpenStudy (anonymous):

The function f(x) is defined for x >= 0 by f(x) = x^a - x^2a a > 0 is a constant How do I solve this for f(x) = 0? where do i find f(x) >0 and f(x) <0

OpenStudy (anonymous):

is it \(x^a-x^{2a}\) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you might try with numbers for \(a\) and see how it looks. for example, if \(a=2\) you get \(x^2-x^4\) that should give you a clue as to how to proceed

OpenStudy (anonymous):

first notice that you can factor, to get \(x^2(1-x^2)\) and in general you can factor and get \(x^a(1-x^a)\)

OpenStudy (anonymous):

How about multiplying the function with \[x ^{-2a}\]

OpenStudy (anonymous):

then you would not have the same function any more, would you?

OpenStudy (anonymous):

That is true

OpenStudy (anonymous):

Okay so now that we have \[x ^{a}(1-x ^{a})\]

OpenStudy (anonymous):

for it to equal 0, either \[x^{a} must be 0\]

OpenStudy (anonymous):

or 1-x^a must be 0

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes i believe so. btw is \(a\) really just some arbitrary constant, or is it a whole number?

OpenStudy (anonymous):

That is what is confusing me as well. On the question sheet is simply says "where a > 0 is a constant"

OpenStudy (anonymous):

not entirely sure how to interpret it

OpenStudy (anonymous):

yeah then i am not either if it was an integer you would have a polynomial, but otherwise you could have something with goofy radicals \[\sqrt[3]{x}(1-\sqrt[3]{x})\] for example

OpenStudy (anonymous):

Why \[\sqrt[3]{x}\]?

OpenStudy (anonymous):

By the way, I would like to tell you that i appreciate you trying to help

OpenStudy (anonymous):

i just picked \(a=\frac{1}{3}\) as an example

OpenStudy (anonymous):

If we assume it is an integer, then would it be right to say \[f(x) \ge 0\] when 0 < x > 1 and \[f(x) \le 0\] when x<0 and x>1

OpenStudy (anonymous):

so it is not clear to me at all what the possible values of \(a\) are perhaps \(a\) could be negative as well right? constant just means some number

OpenStudy (anonymous):

it says a > 0

OpenStudy (anonymous):

sorry if I didnt make that clear

OpenStudy (anonymous):

well still it is not clear. here is what it looks like if \(a=2\) http://www.wolframalpha.com/input/?i=x^2-x^4&dataset= here is what it looks like if \(a=3\) http://www.wolframalpha.com/input/?i=x^3-x^6 try some and see if you can generalize i gotta run good luck

OpenStudy (anonymous):

Thanks! take care

OpenStudy (anonymous):

Oh wow the graphs really helped. If the function is defined for x≥0 then the function is only greater than 0 when x is between 0 and 1 1>x<0